OpenStudy (anonymous):

Solve 2x(x+4)(4x-9)=0.

OpenStudy (anonymous):

Put each bit equal to 0.

OpenStudy (mimi_x3):

x = 0 , -4 , 9/4

OpenStudy (anonymous):

x=0 or x=-4 or x=9/4

OpenStudy (anonymous):

same as prev question: As estudier says, put each "bit" equal to 0 separately :) So, first bit: (2x)=0 x=0 Second bit: (x+4)=0 x=-4 third bit: (4x-9)=0 4x=9 x=9/4